OpenStudy (anonymous):

f"(x) of 3e^(-x)^(2) I dont know how to find this. I am using it so that i can use the trapizodial rule

OpenStudy (turingtest):

is your question \[f(x)=3e^{-x^2}\]find \(f''(x)\) ?

OpenStudy (anonymous):

What is the second derivative?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

can you find the first derivative using the chain rule?

OpenStudy (anonymous):

I thought it would be 3e^(-x^2)*-6x

OpenStudy (turingtest):

if u=-x^2, then du/dx=-2x, so where did the -6x come from?

OpenStudy (anonymous):

i did chain rule on it

OpenStudy (turingtest):

\[u=-x^2\]\[f(x)=3e^{-x^2}=3e^u\]\[f'(x)=3e^u{du\over dx}\]

OpenStudy (turingtest):

what is du/dx ?

OpenStudy (anonymous):

Im really not sure.

OpenStudy (turingtest):

\[u=-x^2\implies{du\over dx}={d\over dx}(-x^2)=?\]